OpenStudy (anonymous):

Jack usually mows his lawn in 3 hours. Marilyn can mow the same yard in 5 hours. How much time would it take for them to mow the lawn together?

OpenStudy (anonymous):

15.8 hours

OpenStudy (anonymous):

simply jack can mow 1/3 of lawn in an hour and marilyn can do 1/5 th of total work they together can mow 1/3+1/5=8/15 th of lawn in an hour and so total time taken to do this togather is 15/8=1.875 h